# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.5.5 Results
In all papers mentioned, the authors claim that the proposed GNN-based models are able to outperform the benchmarks. However, there are a limited but interesting number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs. LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay [art_452].
@@
DST-GCNN [art_126] 2.68 7.20 5.35 3.01 8.50 6.23 3.41 10.30 7.47
DualGraph [art_438] 2.62 7.10 5.36 2.83 7.90 5.89 3.15 9.10 6.67
DyGCN-LSTM [art_113] 2.64 6.35 4.64 3.24 8.04 5.24 3.53 9.30 6.53
@@
MTGNN 2.69 6.86 5.18 3.05 8.19 6.17 3.49 9.87 7.23
MTGODE [art_75] 2.66 6.87 5.10 3.00 8.19 6.05 3.39 9.80 7.05
MVST-GNN [art_109] 2.70 5.19 6.97 3.06 8.12 6.08 3.40 9.60 6.98
PAHNN [art_435] 2.51 6.34 4.82 2.87 7.63 5.73 3.25 9.32 6.67
@@
Clearly, the results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon. However, the multichannel fusion graph neural network (MCFGNN) [art_86] from the “Generic” group consistently exhibits low errors across different time horizons. The MCFGNN model, proposed by Chen et al. in 2022, is composed of three main components: a graph construction layer, a graph convolution layer, and a temporal convolution layer. The input data are first processed by the graph construction layer, which incorporates three channels to capture meta-features, time-varying patterns, and globally stabilized graphs. Spatial and temporal convolutions are then applied in pairs to capture spatio-temporal dependencies in the data. To mitigate the issue of vanishing gradients, residual connections are added to each spatio-temporal block.
@@
MTGNN 1.32 2.77 2.79 1.65 3.69 3.74 1.94 4.53 4.49
MTGODE [art_75] 1.29 2.72 2.73 1.61 3.61 3.66 1.88 4.39 4.31
MVST-GNN [art_109] 1.30 2.73 2.74 1.57 3.56 3.54 1.80 4.19 4.26
PAHNN [art_435] 1.27 2.60 2.65 1.52 3.48 3.54 1.83 4.35 4.27
@@
The results on the the PEMS-BAY dataset vary depending on the metrics and the time horizon. The most accurate models are the adaptive multigraph convolutional networks (AMGCN) [art_444], the multi view spatial-temporal graph neural network (MVST-GNN) [art_109], and the temporal decomposition based dynamic multi granularity graph neural network (TD2MG2NN) [art_422]. The AMGCN model was proposed by Li et al. in 2024, and includes a graph fusion block, an adaptive graph generator block, a stack of spatio-temporal convolution blocks, and an output module. The graph fusion block integrates information from five pre-defined graph structures, based on spatial distance, reachability, correlation, and distributions. Then, an independent adaptive graph generator learns an evolving graph structure to complement the pre-defined graph. Spatio-temporal features are finally extracted through a stack of graph convolutional modules with residual connections, and a time-mixing strategy. The MVST-GNN model, proposed by Li et al. in 2022, includes a multi-view temporal transformer module and a multi-view graph convolution module. The multi-view temporal transformer module is designed to capture dynamic temporal features from recent, daily, and weekly patterns using self-attention and a fusion module. Then, the graph convolution module operates on a self-learned adjacency matrix and passes its output through two fully connected linear layers to generate the final forecast. The TD2MG2NN model, instead, was introduced by Zang et al. in 2024 for evaluating the resilience of road networks in real time during heavy rain events. The process starts by decomposing the time series into seasonal and trend components. These components are processed individually through a season-temporal encoder layer with temporal convolutions, and a trend-temporal encoder layer with regression, respectively. The model then learns two feature matrices, one for daily and one for weekly patterns, which are fed into a dynamic multi-granularity graph convolution module. This module processes the feature matrices in parallel, and then passes them through a spatial decoder layer to produce the final output.
@@
Graph WaveNet 19.85 19.31 32.94
@@
PAHNN [art_435] 14.02 13.19 23.67
STAGCN [art_130] 15.40 14.48 26.23
STCGNN [art_115] 15.81 14.59 27.23
@@
DSTAGNN 19.30 12.70 31.46
DSTGN [art_73] 18.61 12.31 30.79
DSTIGNN [art_74] 18.41 12.45 29.97
@@
SDGL [art_78] 18.65 12.38 31.30
STAGCN [art_130] 19.02 12.46 30.75
STCGNN [art_115] 19.39 12.71 31.17
@@
SDGL [art_78] 14.93 9.61 24.13
SNAS4MTF [art_408] 15.07 9.67 24.00
STAGCN [art_130] 15.36 9.80 24.32
STCGNN [art_115] 15.55 10.02 24.61
@@
As stated above, the results consistently indicate that the periodicity aware spatial-temporal adaptive hypergraph neural network (PAHNN) [art_435] proposed by Zhao et al. in 2024 is the most accurate model. The PAHNN model includes a temporal multi-periodic to time dependencies and periodic features of traffic time series, and a spatial adaptive hypergraph neural network to capture higher-order correlations among nodes.
@@ 5.6 Predictive monitoring
The theme of predictive monitoring is related to sensor-based monitoring of processes, which has become essential in modern industry. The evolution of monitoring systems is concomitant with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [art_9]. Hence, in this context, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system. In the specific, predictive monitoring is a task that involves the continuous observation and analysis of a system in order to forecast its future states and verify whether the predicted outcomes meet certain standards. The predictive monitoring task is important because it can significantly improve efficiency, reliability, and cost of industrial operations. Related data-based approaches include techniques such as anomaly detection, fault diagnosis, and estimation of the equipment’s remaining useful life.
Despite the specificity of the field, 35 out of 281 papers address the topic of predictive monitoring. This includes anomaly detection [art_9, art_10, art_477, art_479, art_481, art_483, art_484, art_489], fault diagnosis [art_53, art_54, art_55, art_56, art_57, art_58, art_192, art_469, art_471, art_474, art_486, art_487, art_491] and RUL estimation [art_152, art_153, art_154, art_155, art_156, art_157, art_158, art_478, art_488]. The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
@@ 5.6.1 Overview
As said, predictive monitoring is a research field which focuses on several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation. Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations from some regular pattern, but are generated by a different underlying mechanism [ref_anomaly_detection_1]. It is a prerequisite part of fault diagnosis [ref_anomaly_detection_2]. Fault diagnosis is the process of determining whether a fault has occurred in a machine, including the identification of the time, location, type, and severity of the fault. It is usually considered to be a classification problem. As for remaining useful life (RUL) estimation, it aims to determine how long a machine will operate before it needs to be repaired or replaced, and it is useful for scheduling maintenance interventions.
The selected papers implement these three techniques, with anomaly detection being less studied than the other two. The main applications developed are industrial (e.g., bearings fault diagnosis [art_54, art_55, art_57, art_155, art_156]) and mechanical (e.g, engines RUL estimation [art_152, art_153, art_154, art_157, art_478, art_488]). There are also some papers dealing with energy related tasks, specifically fault diagnosis of photovoltaic systems [art_53] and of energy networks [art_56, art_487]. They are included in this thematic group because of their specific focus on fault classification rather than on energy-related time series forecasting. It should also be underlined that the majority of these papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis. As for RUL estimation, the quantity to be predicted is the RUL itself. For this reason, we do not make specific reference to forecasting time horizons, unlike in the other subsections, since the goal here is to estimate the RUL rather than to forecast data over future time intervals.
The majority of the collected papers related to this theme were published in 2024 by authors affiliated with Chinese universities. The most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by IEEE, with 6 out of 35 papers. It can be observed that the majority of the articles are published in IEEE journals.
@@ 5.6.2 Datasets
Air Quality China [art_9] https://www.aqistudy.cn/historydata/
@@
Air Quality - TPS [art_9] https://www.kaggle.com/amritpal333/tps-july-2021-original-dataset-clean
C-MAPSS [art_152], [art_153], [art_154], [art_157], [art_478], [art_488] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
@@
CWRU bearing dataset [art_55], [art_192], [art_491] http://csegroups.case.edu/bearingdatacenter/home
Herbert river water level [art_483] https://zenodo.org/records/8053359
Intelligent maintenance system (IMS) [art_475] https://www.nasa.gov/intelligent-systems-division
JNU bearing dataset [art_55] http://mad-net.org:8765/explore.html?id=9&t=0.9355271549540183
@@
NASA battery dataset [art_158] http://ti.arc.nasa.gov/project/prognostic-data-repository
@@
NASA prognostics data [art_58] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/#bearing
@@
N-CMAPSS [art_152], [art_154] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/
MFPT [art_192] https://api.semanticscholar.org/CorpusID:26870683 (source paper)
Mars science laboratory rover (MSL) [art_481], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
NREL PV [art_53] https://www.nrel.gov/
PRONOSTIA [art_155], [art_156] https://hal.science/hal-00719503/document (source paper)
Pooled server metrics (PSM) [art_479], [art_481] https://doi.org/10.1145/3447548.3467174 (source paper)
Server Machine Dataset (SMD) [art_479], [art_484] https://github.com/NetManAIOps/OmniAnomaly
SEU bearing dataset [art_55] https://github.com/cathysiyu/Mechanical-datasets
Secure water treatment (SWaT) [art_477], [art_479], [art_481], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
@@
Soil moisture active passive satellite (SMAP) [art_479], [art_481], [art_484], [art_489] https://doi.org/10.1145/3219819.3219845 (source paper)
@@
Water distribution (WaDi) [art_477], [art_479], [art_484], [art_489] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
Wireless capsule endoscopy images [art_475] https://doi.org/10.1055/s-0043-105488 (source paper)
XJTUGearbox and XJTUSpurgear [art_192], [art_474] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
@@
The two popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA. C-MAPSS describes the deterioration of aircraft engines by recording temperature, pressure, and fan speed during the cruise phase (https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf). N-CMAPSS is an extended version of C-MAPSS, whose records cover climbing, cruising, and descending flight conditions (https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/). Almost no papers use exogenous variables, with the exception of Ref. [art_53], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems.

### v4

@@ 5.5.5 Results
In all papers mentioned, the authors claim that the proposed GNN-based models are able to outperform the benchmarks. However, there are a limited but interesting number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs. LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay.
@@
DST-GCNN [39] 2.68 7.20 5.35 3.01 8.50 6.23 3.41 10.30 7.47
DualGraph [406] 2.62 7.10 5.36 2.83 7.90 5.89 3.15 9.10 6.67
DyGCN-LSTM [198] 2.64 6.35 4.64 3.24 8.04 5.24 3.53 9.30 6.53
@@
MTGNN [424] 2.69 6.86 5.18 3.05 8.19 6.17 3.49 9.87 7.23
MTGODE [177] 2.66 6.87 5.10 3.00 8.19 6.05 3.39 9.80 7.05
MVST-GNN [217] 2.70 5.19 6.97 3.06 8.12 6.08 3.40 9.60 6.98
PAHNN [510] 2.51 6.34 4.82 2.87 7.63 5.73 3.25 9.32 6.67
@@
Clearly, the results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon. However, the multichannel fusion graph neural network (MCFGNN) [52] from the “Generic” group consistently exhibits low errors across different time horizons. The MCFGNN model, proposed by Chen et al. in 2022, is composed of three main components: a graph construction layer, a graph convolution layer, and a temporal convolution layer. The input data are first processed by the graph construction layer, which incorporates three channels to capture meta-features, time-varying patterns, and globally stabilized graphs. Spatial and temporal convolutions are then applied in pairs to capture spatio-temporal dependencies in the data. To mitigate the issue of vanishing gradients, residual connections are added to each spatio-temporal block.
@@
MTGNN [424] 1.32 2.77 2.79 1.65 3.69 3.74 1.94 4.53 4.49
MTGODE [177] 1.29 2.72 2.73 1.61 3.61 3.66 1.88 4.39 4.31
MVST-GNN [217] 1.30 2.73 2.74 1.57 3.56 3.54 1.80 4.19 4.26
PAHNN [510] 1.27 2.60 2.65 1.52 3.48 3.54 1.83 4.35 4.27
@@
The results on the the PEMS-BAY dataset vary depending on the metrics and the time horizon. The most accurate models are the adaptive multigraph convolutional networks (AMGCN) [208], and the spatio-temporal GNN enhanced by a scalable time series pre-training model (STEP) [341]. The AMGCN model was proposed by Li et al. in 2024, and includes a graph fusion block, an adaptive graph generator block, a stack of spatio-temporal convolution blocks, and an output module. The graph fusion block integrates information from five pre-defined graph structures, based on spatial distance, reachability, correlation, and distributions. Then, an independent adaptive graph generator learns an evolving graph structure to complement the pre-defined graph. Spatio-temporal features are finally extracted through a stack of graph convolutional modules with residual connections, and a time-mixing strategy. The STEP model was proposed by Shao et a. in 2021 as a framework to enhance spatio-temporal GNNs through a pre-training approach. Specifically, the pre-training model is designed to learn temporal patterns from long-term historical time series data and generate segment-level representations. These representations, which capture rich contextual information, are then integrated into a spatio-temporal GNN to improve its performance on short-term forecasting tasks. To demonstrate the effectiveness of the framework, the authors use Graph WaveNet as a representative backbone model within STEP.
Other models that, overall, achieve good results are the multi view spatial-temporal graph neural network (MVST-GNN) [217], and the temporal decomposition based dynamic multi granularity graph neural network (TD2MG2NN) [466]. The MVST-GNN model, proposed by Li et al. in 2022, includes a multi-view temporal transformer module and a multi-view graph convolution module. The multi-view temporal transformer module is designed to capture dynamic temporal features from recent, daily, and weekly patterns using self-attention and a fusion module. Then, the graph convolution module operates on a self-learned adjacency matrix and passes its output through two fully connected linear layers to generate the final forecast. The TD2MG2NN model, instead, was introduced by Zang et al. in 2024 for evaluating the resilience of road networks in real time during heavy rain events. The process starts by decomposing the time series into seasonal and trend components. These components are processed individually through a season-temporal encoder layer with temporal convolutions, and a trend-temporal encoder layer with regression, respectively. The model then learns two feature matrices, one for daily and one for weekly patterns, which are fed into a dynamic multi-granularity graph convolution module. This module processes the feature matrices in parallel, and then passes them through a spatial decoder layer to produce the final output.
@@
FourierGNN 17.27 15.88 27.20
Gboot [95] 15.43 14.51 26.42
GMAN 15.52 15.19 26.53
GMSDR [259] 15.78 15.33 26.82
GWN 19.85 19.31 32.94
@@
LSTNet 19.07 17.73 29.67
N-BEATS 18.45 18.35 31.23
PAHNN [510] 14.02 13.19 23.67
PDFormer 21.82 21.47 36.75
PYNet [145] 14.94 14.94 25.27
SFM 17.67 18.33 30.01
SPGCL 23.31 21.88 37.37
STAGCN [112] 15.40 14.48 26.23
STCGNN [438] 15.81 14.59 27.23
StemGNN 14.32 16.24 21.64
@@
DSTAGNN [203] 19.30 12.70 31.46
DSTGN [241] 18.61 12.31 30.79
DSTIGNN [99] 18.41 12.45 29.97
@@
SDGL [243] 18.65 12.38 31.30
STAGCN [112] 19.02 12.46 30.75
STCGNN [438] 19.39 12.71 31.17
StemGNN 21.61 16.10 33.80
STExplainer [369] 18.57 12.13 30.14
@@
SDGL [243] 14.93 9.61 24.13
SNAS4MTF [32] 15.07 9.67 24.00
STAGCN [112] 15.36 9.80 24.32
STCGNN [438] 15.55 10.02 24.61
@@
Z-GCNETs [50] 15.76 10.01 25.11
As stated above, the results almost consistently indicate that the periodicity aware spatial-temporal adaptive hypergraph neural network (PAHNN) [510] proposed by Zhao et al. in 2024 is the most accurate model. The PAHNN model includes a temporal multi-periodic to time dependencies and periodic features of traffic time series, and a spatial adaptive hypergraph neural network to capture higher-order correlations among nodes.
@@ 5.6 Predictive monitoring
The theme of predictive monitoring is related to sensor-based monitoring of processes, which has become essential in modern industry. The evolution of monitoring systems is concomitant with the rapid growth of the Internet of Things (IoT), which aims at a smarter management across various applications [221]. Hence, in this context, multi-sensor systems can be used to detect anomalies in complex scenarios and to monitor the overall status of a system. In the specific, predictive monitoring is a task that involves the continuous observation and analysis of a system in order to forecast its future states and verify whether the predicted outcomes meet certain standards. The predictive monitoring task is important because it can significantly improve efficiency, reliability, and cost of industrial operations. Related data-based approaches include techniques such as anomaly detection, fault diagnosis, and estimation of the equipment’s remaining useful life.
Despite the specificity of the field, 40 out of 366 papers address the topic of predictive monitoring. This includes anomaly detection [221, 226, 515, 249, 114, 21, 512, 136, 128, 272, 283, 239, 204, 139], fault diagnosis [374, 464, 207, 296, 497, 206, 196, 389, 379, 476, 381] and RUL estimation [400, 46, 382, 291, 441, 195, 408, 488, 194]. The use of GNNs in these analyses is justified by the necessity of using a model capable of capturing the spatio-temporal correlations among the many sensors present in the system.
@@ 5.6.1 Overview
As said, predictive monitoring is a research field which focuses on several techniques, such as anomaly detection, fault diagnosis and remaining useful life estimation. Anomaly detection is a technique that aims to identify abnormal data that are not due to random deviations from some regular pattern, but are generated by a different underlying mechanism [131]. It is a prerequisite part of fault diagnosis [172]. Fault diagnosis is the process of determining whether a fault has occurred in a machine, including the identification of the time, location, type, and severity of the fault. It is usually considered to be a classification problem. As for remaining useful life (RUL) estimation, it aims to determine how long a machine will operate before it needs to be repaired or replaced, and it is useful for scheduling maintenance interventions.
The selected papers implement these three techniques, with anomaly detection being less studied than the other two. The main applications developed are industrial (e.g., bearings fault diagnosis [464, 207, 296, 291, 441]) and mechanical (e.g, engines RUL estimation [400, 46, 382, 195, 488, 194]). There are also some papers dealing with energy related tasks, specifically fault diagnosis and degradation analysis of photovoltaic systems [374, 79] and of energy networks [476]. They are included in this thematic group because of their specific focus on fault classification rather than on energy-related time series forecasting. It should also be underlined that the majority of these papers aim at solving a classification task, which is typical of anomaly detection and fault diagnosis. As for RUL estimation, the quantity to be predicted is the RUL itself. For this reason, we do not make specific reference to forecasting time horizons, unlike in the other subsections, since the goal here is to estimate the RUL rather than to forecast data over future time intervals.
The majority of the collected papers related to this theme were published in 2024, and the most popular journal for these topics is IEEE Transactions on Instrumentation and Measurement by IEEE, with 6 out of 35 papers. It can be observed that the majority of the articles are published in IEEE journals. As for conference venues, 4 out of 7 conference papers were presented at CIKM.
@@ 5.6.2 Datasets
Air quality China [221] https://www.aqistudy.cn/historydata/
@@
Air quality - TPS [221] https://www.kaggle.com/amritpal333/tps-july-2021-original-dataset-clean
C-MAPSS [400], [46], [382], [195], [488], [194] https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf
@@
CWRU bearing dataset [207], [206], [381] http://csegroups.case.edu/bearingdatacenter/home
HAI security dataset [128] https://github.com/icsdataset/hai
Herbert river water level [21] https://zenodo.org/records/8053359
Intelligent maintenance system (IMS) [96] https://www.nasa.gov/intelligent-systems-division
JNU bearing dataset [207] http://mad-net.org:8765/explore.html?id=9&t=0.9355271549540183
MFPT [206] https://api.semanticscholar.org/CorpusID:26870683 (source paper)
Mars science laboratory rover (MSL) [114], [136], [272], [204] https://doi.org/10.1145/3219819.3219845 (source paper)
@@
NASA battery dataset [408] http://ti.arc.nasa.gov/project/prognostic-data-repository
@@
NASA prognostics data [497] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/#bearing
@@
N-CMAPSS [400], [382] https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/
NREL PV [374] https://www.nrel.gov/
PRONOSTIA [291], [441] https://hal.science/hal-00719503/document (source paper)
Pooled server metrics (PSM) [249], [114] https://doi.org/10.1145/3447548.3467174 (source paper)
PV degradation [79] https://github.com/Yangxin666/ST-GTrend/tree/main/Datasets
SEU bearing dataset [207] https://github.com/cathysiyu/Mechanical-datasets
@@
Soil moisture active passive satellite (SMAP) [249], [114], [512], [136], [204] https://doi.org/10.1145/3219819.3219845 (source paper)
Server machine dataset (SMD) [249], [512], [272] https://github.com/NetManAIOps/OmniAnomaly
@@
Secure water treatment (SWaT) [515], [249], [114], [512], [136], [128], [272], [283], [239], [204] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
Water distribution (WaDi) [515], [249], [512], [136], [128], [272], [283], [239], [204] https://itrust.sutd.edu.sg/itrust-labs_datasets/dataset_info/
Wireless capsule endoscopy images [96] https://doi.org/10.1055/s-0043-105488 (source paper)
XJTUGearbox and XJTUSpurgear [206], [389] http://dx.doi.org/10.1016/j.ymssp.2021.108653 (source paper)
@@
The two popular datasets are C-MAPSS and N-CMAPSS, two well-known datasets for RUL prediction provided by NASA. C-MAPSS describes the deterioration of aircraft engines by recording temperature, pressure, and fan speed during the cruise phase (https://ntrs.nasa.gov/api/citations/20070034949/downloads/20070034949.pdf). N-CMAPSS is an extended version of C-MAPSS, whose records cover climbing, cruising, and descending flight conditions (https://ti.arc.nasa.gov/tech/dash/groups/pcoe/prognostic-data-repository/). Almost no papers use exogenous variables, with the exception of Ref. [374], which evaluates the impact of meteorological variables on fault diagnosis of photovoltaic systems. Other popular datasets are SWaT and WaDi. The secure water treatment dataset (SWaT) is a cybersecurity dataset that includes data collected from 51 sensors and actuators over a continuous period of 11 days. During this time, the system operated under normal conditions for 7 days and was subjected to attack scenarios for 4 days. The dataset contains both network traffic and physical process data recorded from all devices. The water distribution (WaDi) dataset is an extension of SWaT involving a larger setup of 127 sensors and actuators monitored for 16 consecutive days.